Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 63 points. They won 12 more games than they tied. How many wins and how many ties did the team have?
Answer by jorel1380(3719)

2x+y=63
x=y+12
2(y+12)+y=63
3y+24=63
3y=39
y=13
x=25
The hockey team won 25 games and tied 13..